OpenStudy (lilai3):

Glen received 6 birthday cards. If he is equally likely to read the cards in any order, what is the probability he reads the card from his parents and the card from his sister before the other cards?

OpenStudy (experimentx):

question unclear ... if you mean in order than, it's 1/6x1/5

OpenStudy (lilai3):

i typed dat straight out of my hmwk...-__-

OpenStudy (eyust707):

your homework has just slightly better grammar

OpenStudy (lilai3):

? wut u mean by dat... eyust707?

OpenStudy (eyust707):

hehe im just teasing you.

OpenStudy (lilai3):

tsh..... :)

OpenStudy (experimentx):

well, i think you should really be clear about the situation presented by the question

OpenStudy (lilai3):

is it 1/6? cuz den it says here dat 'he reads the cards from his parents and the cards from his sister b4 da other cards?' wait..is dat suppose 2 b 1 card altogether or 2..or 3?

OpenStudy (experimentx):

then it's 1/30 ... fist one from parent THEN from sister before others means he reads in specific order

OpenStudy (lilai3):

u're absolute ( from your instinicts) that the asnwer is 1/20... right?

OpenStudy (anonymous):

\[\left(\frac{2}{6}\right)\left(\frac{1}{5}\right)\text{=}\frac{1}{15}\]The probability of selecting a card either from the parents or the sister is 2/6. The probability of selecting the other card from the sister or from the parents is 1/5.

OpenStudy (lilai3):

thank you very much!

OpenStudy (anonymous):

Your are welcome. Thank you for the medal.

OpenStudy (lilai3):

yw :)

OpenStudy (lilai3):

and good nite, amigos (btw, don't think i'm spanish... cuz i'm not!)